OpenStudy (anonymous):

solve inequality and graph solution -3e - 10 less than or equal to -4 -3e - 10 +10 less than or equal to -4 +10 -3e less than or equal to 5 -3e/3 less than or equal to 5/3 e less than or equal to 1.6666

OpenStudy (anonymous):

-4+10 should be 6 not 5

OpenStudy (anonymous):

Then when you divide by -3, it flips the inequality sign

OpenStudy (anonymous):

\[-3e \le 6\] \[e \ge -2\]

OpenStudy (anonymous):

ok I got it so I would have an open circle on -2 and it would go towards the positive side

OpenStudy (anonymous):

Shaded circle since its less than or EQUAL to, and yes shading to the right

OpenStudy (anonymous):

so you would flip the sign if the other side is positive correct

OpenStudy (anonymous):

You flip the sign whenever you divide or multiply by a negative number we divided by -3 so we flipped the sign

OpenStudy (anonymous):

so did I do this one right 15 less than or equal to 3 -4s 15-3 less than or equal to 3-3 -4s 12 less than or equal to -4s 12/-4 great than or equal to -4s/-4 -3 great than or equal to s since this is not true it would be an open circle towards the negative

OpenStudy (anonymous):

You're right with \[-3 \ge s\] but it would be shaded circle on -3, shading to the left

OpenStudy (anonymous):

oh is it because the -3(-4) would be 12

OpenStudy (anonymous):

No, its because s is less than or equal to -3 so that means -3 and everything less than it is included

OpenStudy (anonymous):

ok thanks